OpenStudy (4everaddicted2anime):

Medal and Fan. Find the missing measure for a right circular cone given the following information. Find V if r = 2 and L.A. = 12pi.

OpenStudy (4everaddicted2anime):

Answer choices:\[A=(16/3)\sqrt{2\pi}\]\[B=16\sqrt{2\pi}\]\[C=4\sqrt{38\pi}\]

OpenStudy (4everaddicted2anime):

@anthony1914

OpenStudy (4everaddicted2anime):

The height is the square root of 32

rishavraj (rishavraj):

my bad ... i mentiond for cylinder. ...

rishavraj (rishavraj):

so for cone ...u got the la and r \[2\pi \sqrt{h^2 + 4} = 12\]

rishavraj (rishavraj):

get h from above equation...and volume = \[\frac{ \pi r^2h }{ 3 }\]

rishavraj (rishavraj):

it 12 *pi

OpenStudy (4everaddicted2anime):

Did I get h right? If not, how do I solve for h underneath the square root sign?

rishavraj (rishavraj):

\[\sqrt32 = 4\sqrt2\]

rishavraj (rishavraj):

@4everaddicted2anime there ??? and the first option is it \[\frac{ 16 }{ 3 }\sqrt{2} ~\pi ~~or~~\frac{ 16 }{ 3 }\sqrt{2\pi}\]

OpenStudy (4everaddicted2anime):

Thank you